You are presented with the following equation: (1+r) = (1+i ) / (1 +π).

Expand the expression to solve for i. What assumption is required for this expression to be equal to the Fischer equation?
Answer #1

Cross multiply will result in

1 + π + r + πr = 1 + i

This gives i = π + r + πr

This is the expression to solve the nominal interest rate

We know that fischer equation is i = r + π. The above expression can be converted into Fischer equation if we assume that πr is approximately 0.